OpenStudy (anonymous):

Choose the correct first step(s) to solve the system using elimination. 4x + 7y = 6 and 6x + y = 2

OpenStudy (anonymous):

Multiply 6x+y=2 by -7, but there are many choices possible. Is this multiple choice?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

Listing the choices would help figure it out.

OpenStudy (anonymous):

sure here.7(6x + y = 2) → -42x - 7y = -14 4x + 7y = 6x + y y = -6x + 20 → 4x + 7(-6x + 20) = 6 (4x + 7y) + (6x + y) = 8 (4x + 7y) - (6x + y) = 4

OpenStudy (anonymous):

-42x-7y=-14

OpenStudy (anonymous):

multiplying by -7

OpenStudy (anonymous):

i thought that one also. between and third one

OpenStudy (anonymous):

The third one wouldn't eliminate any variables though.

OpenStudy (anonymous):

aha ok. i'll give you a medal when i find out if its correct or not :D